OpenStudy (anonymous):

how do i find the zeros of x^4-5x^3+8x^2-10x+12 ? i already found x=2, but im not sure how to find the rest bc i used synth. division but didnt find any more real zeros and i cant remember how to find imaginary zeros. help!

OpenStudy (dumbcow):

ok after factoring out the (x-2) you get x^3-3x^2+2x-6 factor this by grouping x^2(x-3) + 2(x-3) (x^2+2)(x-3) x-3=0 --> x=3 x^2+2 =0 x^2=-2 x=+-sqrt(-2) x = +-sqrt(2)i

OpenStudy (anonymous):

thank you!!!

OpenStudy (dumbcow):

your welcome :)